Calculate the coumpound interest.
A(t)=P(1+r/n)^nt
How much would $200 invested at 4% interest compounded monthly be worth after 30 years?

Answer by checkley77(12844)

A(30)=200(1+.04/12)^12*30
A(30)=200(1+.003333)^360
A(30)=200(1.003333)^360
A(30)=200*3.3131
A(30)=$662.62 ans.